The Duval Company was incorporated with 8% preferred capital stock of $600,000 and common stock of $2,200,000. The par value of the preferred stock was $100, and the par value of the common stock was $20.

Last year, dividends were declared by the Duval Company, which had earnings totaling $422,000.

a. What was the total amount of the preferred stock dividend?

$

b. What amount would have been paid on each share of common stock if all the earnings had been distributed?

$ per share
Answer #1

a) Total amount of preferred stock dividend = 600000*8% = 48000

b) Common stock dividend per share = (422000-48000)/110000 = 3.40 per share